OpenStudy (anonymous):

Convert to decimal notation. 10^-3

OpenStudy (anonymous):

\[10^{-3}=\frac{1}{10^3}=\frac{1}{1000}=0.001\]

OpenStudy (anonymous):

really just put down a "1" and move the decimal 3 places left